# Mamo: a Mathematical Modeling Benchmark with Solvers

### v1

@@ 6.2 Benchmarking Results
Our evaluation includes both closed-source LLMs and open-source LLMs. The closed-source commercial LLMs include the GPT-4 series OpenAI (2023), the Claude series  [^4] https://www.anthropic.com/news/claude-3-family, and the Gemini series  [^5] https://deepmind.google/technologies/gemini/. The open-source LLMs include DeepSeek-v2 DeepSeek-AI (2024) and DeepSeek-coder Guo et al. (2024). Table 2 presents the evaluation results of different language models on the Mamo Benchmark. The performance of these models is assessed both in their original form and when enhanced with the language model itself to rectify syntax errors (indicated by the dagger symbol ). Table 3 displays the performance results when GPT-4-0613 is selected as the code modifier.
Models ODE Optimization
First order Second order System All Easy LP Complex LP
GPT-4o †† 69.90% 74.55% 42.50% 68.21% 87.42% 29.86%
GPT-4-turbo †† 67.85% 70.00% 42.50% 65.99% 88.19% 28.44%
GPT-4 †† 61.73% 67.27% 45.00% 61.56% 86.96% 24.64%
GPT-3.5-turbo †† 47.96% 60.91% 20.00% 48.84% 85.43% 10.43%
Claude-3-Sonnet †† 58.16% 66.36% 35.00% 59.09% 85.12% 28.44%
Claude-3-Haiku †† 53.06% 65.45% 25.00% 53.76% 86.81% 21.33%
Gemini-1-pro †† 43.37% 54.55% 15.00% 43.64% 80.37% 16.27%
Gemini-1.5-pro †† 67.35 % 67.27% 52.50% 65.61% 85.28% 34.12%
DeepSeek-v2 †† 56.12% 68.18% 32.50% 57.23% 86.81% 6.16%
DeepSeek-coder †† 33.67% 49.09% 17.50% 36.71% 78.83% 18.96%
@@
The evaluation results on the Mamo Benchmark reveal that GPT-4o stands out among the models in both ODE and LP tasks. Specifically, GPT-4o achieves the highest performance in ODE tasks with 69.90% in first-order and 74.55% in second-order equations, and also excels in LP tasks with 87.42% in easy LP and 32.23% in complex LP scenarios when using the LLM to rectify syntax errors. Additionally, DeepSeek-v2, as an open-source LLM, demonstrates significant capability with 48.27% overall in ODE tasks and 86.66% in easy LP tasks when enhanced with syntax correction, highlighting its potential despite being open-source.
The inclusion of the LLM itself as code modifier to rectify syntax errors results in significant improvements across various models. For instance, GPT-3.5 Turbo’s performance in ODE tasks increased from 15.90% to 26.30% overall and Claude-3-Haiku shows a remarkable improvement in ODE tasks from 45.66% to 52.31%. Similarly, DeepSeek-v2, as an open-source LLM, demonstrates significant capability with its overall ODE performance increasing from 41.33% to 48.27% when enhanced with syntax correction.
@@
Selecting GPT-4 as a code modifier generally results in better performance than using the LLM itself for self-improvement. For instance, DeepSeek-v2’s accuracy in ODE tasks improves from 48.27% with self-improvement to 57.23% when GPT-4 is used as the code modifier. This pattern indicates that GPT-4 is a more effective code modifier, enhancing the accuracy and performance of various models more reliably than self-improvement.
6.3 On the format errors and code modifiers
Models ODE Optimization
format raw self-improved GPT-4-improved raw self-improved GPT-4-improved
GPT-4o 96.82% 99.42% 99.42% 92.93% 97.91% 97.91%
GPT-4-turbo 96.82% 99.13% 98.84% 93.86% 97.80% 97.56%
GPT-4 94.51% 97.69% 97.69% 90.96% 95.13% 95.13%
GPT-3.5-turbo 32.37% 60.40% 94.22% 87.95% 89.80% 96.76%
Claude-3-Sonnet 84.68% 93.06% 97.40% 91.89% 93.97% 96.76%
Claude-3-Haiku 78.03% 92.77% 98.27% 82.61% 92.58% 96.29%
Gemini-1.0-pro 56.36% 69.94% 88.73% 88.06% 89.11% 94.21%
Gemini-1.5-pro 93.64% 98.27% 98.84% 95.48% - 96.87%
deepseek-v2 70.81% 83.81% 96.24% 90.27% 92.00% 94.90%
deepseek-coder 88.44% 93.64% 95.95% 92.58% 96.29% 98.84%
Table 4: Correct format rate on raw, self-improved, and GPT4-improved format.
The data presented in Table 4 provides a comparative analysis of the performance (rate of correct format) of various models using different improvement methods. The table considers three scenarios: no improvement, self-improvement, and improvement provided by GPT-4.
The GPT-3.5-turbo model shows the most dramatic improvement when enhanced by GPT-4, jumping from a 32.37% success rate to 94.22%. Interestingly, Claude-3 models (Sonnet, and Haiku) and DeepSeek models (DeepSeek-v2, DeepSeek-coder) also exhibit high improvement rates when enhanced by GPT-4, highlighting the compatibility of GPT-4’s improvement methods with other models. In the realm of Optimization tasks, the GPT-3.5-turbo model again shows a significant improvement when enhanced by GPT-4, jumping from a 87.95% success rate to 96.76%.
The improvement from GPT-4 is generally superior to both self-improvement and no improvement. This indicates that GPT-4 can serve as a more effective and fair code modifier. By using GPT-4 for syntax error correction, we can better examine the modeling ability of different LLMs without the confounding factor of formatting errors. Notably, in ODE tasks, the correct format rate for all models, except one, is generally greater than 94%, with the lowest being 88.73%. In Optimization tasks, the correct format rate is even higher, with the lowest rate being 94.21%.
7 Conclusion and Future Directions
In this article, we introduce Mamo, a novel benchmark that utilizes solvers to evaluate the mathematical modeling capabilities of large language models (LLMs). We propose a shift in the evaluation paradigm from traditional outcome-oriented evaluation to process-oriented evaluation. This shift allows us to scrutinize the modeling process undertaken by LLMs, providing a more detailed and comprehensive assessment of their mathematical problem-solving abilities. Our evaluation focuses more on the mathematical modeling capabilities of the models, leaving the process of solving abstract mathematical problems to the solvers, thereby avoiding the impact of errors such as computational capabilities of the models on their actual modeling abilities.
Our benchmark represents an important step in understanding the fundamental modeling capabilities of LLMs. By focusing on the modeling process rather than the correctness of the final solution, we are able to gain valuable insights into the strengths and limitations of current LLMs. This approach provides new directions for future research, emphasizing the importance of the modeling process in solving mathematical problems.
The introduction of Mamo paves new avenues for future research. The benchmark can be expanded to include a wider variety of mathematical problems and solvers, broadening its applicability and providing a more comprehensive evaluation of the mathematical modeling capabilities of LLMs. Insights gained from the benchmark can be used to guide the development of future LLMs. By understanding the strengths and weaknesses of current models in solving mathematical problems, researchers can design new models that better handle these tasks. We will extend our model list to other models (Llama2 series, Mistral series etc) in our further study.
Limitations
Ethics Statement:
Appendix A Cross Review
@@
We have also conducted the cross-review process. See the details in Appendix.  A.1 and the results on Appendix.  A.1.1.
In the optimization section, particularly with easy LP part, out of 688 entries, 12 were found to be completely invalid (not meeting the criteria), 26 had incorrect answers, and 6 had misleading descriptions. All invalid entries were filtered out, and the remaining issues were corrected manually. Among them, 8 of them were corrected and others (36 questions) were filtered out. In complex LP part, for the original 211 questions, 40 were found synthetized by wrong mathematical models. All of them were corrected.
A.1 Cross-Review Process
Furthermore, adhering to the scaling law evident in the dataset’s distribution, we selected 50 questions from the ODE parts in our dataset for a deeper analysis. Four independent reviewers (see their qualification in Section B) were tasked with assessing 50 questions each. The responses could either be numeric or labeled as “error” if a question was deemed problematic. This section outlines the metrics used to evaluate the reviewers’ responses against pre-defined correct answers, which also include the “error” label for any flawed questions. We use the same metric as described in Subsection 6.1.2. Here are our results:
A.1.1 Results
@@
Appendix B Annotator Qualifications
@@
B.1 Qualification of Reviewers
We select the reviewers mentioned in Appendix  A.1 based on the following criteria:
@@ B.1 Qualification of Reviewers
Appendix C Few-shot Experiment
To test the sensitivity of different models to the number of prompts, we conducted a sensitivity analysis using the scaling law to examine the effect of the number of shots on model performance. In our analysis, we tested the performances of 11 models on ODE problems using prompts with 0 shot, 1 shot, 3 shots, 5 shots, and 10 shots. The results, shown in Figure  3, illustrate the ODE problem-solving accuracy of different models under varying numbers of shots. Similarly, we performed sensitivity tests on optimization problems, assessing the accuracy of 10 models with prompts of 0 shot, 1 shot, 3 shots, and 5 shots. The results are depicted in Figure  4.

### v2

@@ 6.2 Benchmarking Results
Our evaluation  [^6] See the setting in Appendix N includes both closed-source LLMs and open-source LLMs. The closed-source commercial LLMs include the GPT-4 series OpenAI (2023), the Claude series  [^7] https://www.anthropic.com/news/claude-3-family, and the Gemini series  [^8] https://deepmind.google/technologies/gemini/. The open-source LLMs include DeepSeek series by  DeepSeek-AI (2024),  Guo et al. (2024) and  Shao et al. (2024); Llama-3 models  [^9] https://github.com/meta-llama/llama3; Qwen-1.5 (Bai et al.) and Mixtral (Jiang et al.). Table 3 presents the evaluation results of different language models on the Mamo Benchmark. The performance of these models is assessed both in their original form and when enhanced with the language model itself to rectify syntax errors (the improvement is indicated by the subscript). Table 5 in Appendix A shows results when GPT-4-0613 is selected as the code modifier.
The evaluation results on the Mamo Benchmark reveal that GPT-4o stands out among the models in both ODE and LP tasks. Specifically, GPT-4o achieves the highest performance in ODE tasks with 68.37% in first-order and 37.27% in second-order equations, and also excels in LP tasks with 87.42% in easy LP and 31.28% in complex LP scenarios when using itself as code modifier. Additionally, DeepSeek-v2, as an open-source LLM, demonstrates significant capability, achieving a 57.57% success rate across all tasks when enhanced with syntax correction by itself, underscoring its potential despite its open-source nature. Moreover, Mixtral-8x22B also exhibits competitive performance relative to DeepSeek-v2. After fix syntax error by itself, it achieves a remarkable score of 21.33% on the complex LP task, showcasing its potential in modeling complex optimization problems.
The inclusion of the LLM itself as code modifier to rectify syntax errors results in significant improvements across various models. For instance, GPT-3.5-turbo’s performance in ODE tasks increased from 12.72% to 18.21% overall and Claude-3-Haiku shows a remarkable improvement in ODE tasks from 32.95% to 38.73%. Similarly, DeepSeek-v2, as an open-source LLM, demonstrates significant capability with its overall ODE performance increasing from 32.08% to 35.55% when enhanced with syntax correction.
6.3 On the Format Errors and Code Modifiers
Table 4: Execution Success Rates for Raw, Self-Modified, and GPT-4 Modified Methods, denoting scenarios without code modifiers, utilizing the LLM as a code modifier, and employing GPT-4-0613 as the code modifier. This table highlights the success rates of file execution (correct format) across these three scenarios, differing from the Table 3 which shows the correctness rate.
Models ODE (%) Optimization (%)
raw self-modified GPT-4-modified raw self-modified GPT-4-modified
GPT-4o 96.82 99.42 99.42 92.93 97.91 97.91
GPT-4-turbo 96.82 99.13 98.84 93.86 97.80 97.56
GPT-4 94.51 97.69 97.69 90.96 95.13 95.13
GPT-3.5-turbo 37.28 65.03 93.93 87.95 89.80 96.76
Claude-3-Sonnet 84.68 93.06 97.40 91.89 93.97 96.76
Claude-3-Haiku 78.03 92.77 98.27 82.61 92.58 96.29
Claude-3-Opus 88.73 96.82 97.98 90.38 93.40 95.94
Gemini-1.0-pro 56.36 69.94 88.73 88.06 89.11 94.21
Gemini-1.5-pro 93.64 98.27 98.84 95.48 98.73 96.87
DeepSeek-v2 70.81 83.82 96.24 90.27 92.00 94.90
DeepSeek-coder 88.44 93.64 95.95 92.58 96.29 98.84
DeepSeek-math-7B-base 28.90 29.77 79.77 55.50 58.40 75.67
DeepSeek-math-7B-rl 5.20 6.64 80.64 19.93 21.55 55.97
Llama-3-8B 20.23 26.01 68.21 75.09 76.36 79.14
Llama-3-70B 24.56 30.92 77.46 77.64 79.49 89.46
Qwen-1.5-72B-Chat 55.49 89.60 95.38 77.40 92.24 95.37
Qwen-1.5-110B-Chat 51.45 63.29 95.09 71.26 84.01 90.15
Mixtral-8x7B-instruct 50.87 69.08 91.62 84.13 87.83 96.29
Mixtral-8x22B-instruct 80.92 88.73 96.24 87.72 93.51 96.29
The data presented in Table 4 provides a comparative analysis of the performance (rate of correct format) of various models using different modification methods. The table considers three scenarios: no modification, self-modification, and modification provided by GPT-4-0613.
The GPT-3.5-turbo model shows the most dramatic improvement when modified by GPT-4, jumping from a 37.28% success rate to 93.93%. Interestingly, Claude-3 models (Sonnet and Haiku) and DeepSeek models (DeepSeek-v2, DeepSeek-coder) also exhibit high improvement rates when enhanced by GPT-4-0613, highlighting the compatibility of GPT-4’s modification methods with other models. In the realm of Optimization tasks, the GPT-3.5-turbo model again shows a significant improvement when enhanced by GPT-4, jumping from a 87.95% success rate to 96.76%.
The modification from GPT-4 is generally superior to both self-modification and no modification. This indicates that GPT-4 can serve as a more effective and fair code modifier. By using GPT-4 for syntax error correction, we can better examine the modeling ability of different LLMs without the confounding factor of formatting errors. Table 5 in Appendix A shows the results of selecting GPT-4-0613 as code modifier.
6.4 Analysis
Given the results in Section 6.26.3 and Appendix A, several observations emerge from the benchmark results. For open-source models, larger models generally perform better, as evidenced by comparisons between Llama-3-70B and Llama-3-8B, Mixtral-8x22B-instruct and Mixtral-8x7B-instruct, with the larger models consistently achieving higher scores.
From Table 4, it is apparent that some models struggle with adhering to the syntax of the solver, such as GPT-3.5-turbo and certain open-source models. Some models like Llama-3-8B may not closely follow the instructions, even worse than the syntax error.These models tend to show significant improvements in both correct rates (as indicated by the subscript in Table 3) and execution success rates after the application of code modifiers, particularly when their initial performance without modifiers is low. Notably, in the ODE part, GPT-3.5-turbo achieves a remarkable 28.03% increase in execution success rate by self-modified, accompanied by substantial improvements in ODE performance. In Optimization part, Qwen-1.5-72B-Chat follows the same trends. When comparing the results from Table 3 and Table 5, it becomes clear that the improvements are even more pronounced for open-source models when utilizing GPT-4 modifications rather than self-modifications. See Appendix E for the analysis of typical modeling errors and case study,
7 Conclusion and Future Directions
In conclusion, Mamo’s innovative benchmark shifts the evaluation focus of LLMs towards the modeling process, offering a nuanced assessment of their mathematical problem-solving capabilities. By isolating modeling from solving through the use of solvers, Mamo provides a comprehensive understanding of LLMs’ strengths and limitations in mathematical modeling. This approach not only enhances our insights into the fundamental modeling abilities of LLMs but also guides future research by highlighting the importance of the modeling process in solving mathematical problems.
Appendix A GPT-4-0613 as Code Modifier
@@
Models ODE Optimization Overall (%)
First Second System Easy Complex
order (%) order (%) (%) LP (%) LP (%)
GPT-4o †† 68.37 37.27 40.00 87.42 29.38 68.07
GPT-4-turbo †† 65.82 33.64 40.00 88.19 27.96 67.58
GPT-4 †† 60.20 29.09 42.50 86.96 24.17 64.93
GPT-3.5-turbo †† 44.90 20.00 20.00 85.43 9.48 57.48
Claude-3-Opus †† 59.18 36.36 47.50 84.20 14.22 62.37
Claude-3-Sonnet †† 56.63 26.36 32.50 85.12 27.96 63.44
Claude-3-Haiku †† 51.53 25.45 22.50 86.81 18.96 61.54
Gemini-1-pro †† 40.31 18.18 10.00 80.21 15.64 54.51
Gemini-1.5-pro †† 65.82 30.91 50.00 85.28 34.12 67.08
DeepSeek-v2 †† 54.08 30.91 30.00 86.81 3.79 60.05
DeepSeek-coder †† 30.10 11.82 15.00 78.83 18.96 52.27
DeepSeek-math-7b-base †† 13.78 9.09 5.00 59.05 9.48 36.72
DeepSeek-math-7b-rl †† 13.78 7.27 7.50 47.70 1.90 29.20
Llama-3-8b †† 11.22 0.91 0.00 63.19 8.53 37.47
Llama-3-70b †† 14.80 7.27 10.00 70.71 9.48 43.18
Qwen-1.5-72b-Chat †† 16.84 13.01 5.45 78.53 6.64 47.80
Qwen-1.5-110b-Chat †† 40.82 13.64 20.00 78.83 14.22 53.51
Mixtral-8x7b-instruct †† 32.14 11.82 20.00 75.46 9.95 49.38
Mixtral-8x22b-instruct †† 55.10 17.27 22.50 85.43 22.27 61.21
@@
Selecting GPT-4 as a code modifier generally results in better performance than using the LLM itself for self-improvement. For instance, DeepSeek-v2’s accuracy in ODE tasks improves from 32.08% with self-modification to 43.93% when GPT-4 is used as the code modifier. This pattern indicates that GPT-4 is a more effective code modifier, enhancing the accuracy and performance of various models more reliably than self-improvement.
Appendix B Cross Review
@@
We have also conducted the cross-review process. See the details in Appendix B.1 and the results on Appendix B.1.1.
In the optimization section, particularly with easy LP part, out of 688 entries, 12 were found to be completely invalid (not meeting the criteria), 26 had incorrect answers, and 6 had misleading descriptions. All invalid entries were filtered out, and the remaining issues were corrected manually. Among them, 8 of them were corrected and others (36 questions) were filtered out. In complex LP part, for the original 211 questions, 40 were found formulated by wrong mathematical models. All of them were corrected.
B.1 Cross-Review Process
Furthermore, adhering to the scaling law evident in the dataset’s distribution, we selected 50 questions from the ODE parts in our dataset for a deeper analysis. Four independent reviewers (see their qualification in Section C) were tasked with assessing 50 questions each   [^10] The hourly wage pay is 100 RMB/h, the total amount spent on the participant compensation is 2400 RMB, approximately 331 dollars. The responses could either be numeric or labeled as “error” if a question was deemed problematic. This section outlines the metrics used to evaluate the reviewers’ responses against pre-defined correct answers, which also include the “error” label for any flawed questions. We use the same metric as described in Subsection 6.1.2. Here are our results:
B.1.1 Results
@@
Appendix C Annotator Qualifications
@@
C.1 Qualification of Reviewers
We select the reviewers mentioned in Appendix B.1 based on the following criteria:
@@ C.1 Qualification of Reviewers
Appendix D Metric
The comparison process is as the following:
$\text{Correctness}=\begin{cases}1&\text{if }\left|\frac{O-A}{A}\right|\leq 1\times 10^{-4}\text{~{}{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0} (1)} or }|O-A|<1\times min\{10^{-n},10^{-2}\}\text{~{}{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0} (2)},}\\ 0&\text{otherwise.}\end{cases}$
Our consideration: (1) The difference is much less than the Answers, so can be omitted.
(2) The only difference lies in the last few digits of the required significant figure of the answer. By selecting different thresholds $\xi\times min\{10^{-n},10^{-2}\}$, we can adjust the tolerance for the discrepancy, ensuring that $|O-A|<\xi\times min\{10^{-n},10^{-2}\}$. We have explored the relationship between various thresholds and the accuracy of GPT-4o.
Table 6: Accuracy at various thresholds for different categories(%), where $\xi_{i}=i$
$\xi$ 0 1 2 3 4 5 6 7 8 9
ODE 48.55 50.87 54.62 56.36 57.51 58.96 59.83 60.12 60.69 61.85
Easy LP 87.27 87.27 87.27 87.27 87.27 87.27 87.27 87.27 87.27 87.27
Complex LP 22.75 22.75 22.75 22.75 22.75 22.75 22.75 22.75 22.75 22.75
Figure 2: Accuracy of different thresholds in different categories
Figure 2 plots the results. In our metric, we choose strict $\xi=1$.
Appendix E Error Analysis
In this section, we will analyze some typical errors.
E.1 Deviation from Instruction
It is common for some models that show a low capability of few-shot learning, especially for small models. For example, in a case Llama-3-8B output ‘I wish you good luck’ when prompting with some few-shot demonstrations in mathematical modeling.
E.2 Syntax Error
In ODE parts, the syntax errors is obviously the syntax errors in the reponsed python code. In Optimization part, the .lp file require more strict syntax than the universal optimization formulation. The Figure 4 and 4 show the typical syntax errors in .lp format.
Figure 3: The syntax error is caused by the constraint, in .lp format, the variables should be placed on the left hand side of the inequality
Figure 4: Same type of syntax error as Figure 4
E.3 Modeling Error
One of the most common errors in modeling is the lack of recognition of decision variable types. The types of decision variables in optimization problems are divided into discrete (sometimes integer, sometimes binary) and continuous. Figure 5 is an example.
Figure 5: In this TSP problem, the decision variable should be binary, instead of "General", which means “integer”
Appendix F Scope of Mamo
Table 7: The taxonomy of mathematical modeling. In the Mamo benchmark, we only include the categories where sophisticated solvers are available.
Category In Mamo
Changes and Differences
Proportionality and Geometric Similarity
Optimization Problem Modeling
Differential Equations
Probabilistic Modeling
The scope of Mamo is chosen under the consideration in Section 4.2. As Table 7 shows, Mamo focus on the ODE and Optimization areas.
Appendix G Examples of Different ${\widehat{M}}$
Different models will typically lead to different final answers after solving by the solvers. For example, in ODE part:
${\widehat{M}}_{1}:y^{\prime}+y=0$
and
${\widehat{M}}_{2}:y^{\prime}+2y=0$
Suppose we ask for $y(1)$, then we have ${\widehat{A}}_{1}=e^{-1}$ and ${\widehat{A}}_{1}=e^{-2}$,they have different values.
Then for the optimization,
${\widehat{M}}_{1}$:
$\max_{x,y}$ $x+y$
subject to $x+y\leq 20,$
$x-y\geq 0,$
$0\leq x\leq 10,$
$0\leq y\leq 5,$
$x,y\in\mathbb{Z}.$
${\widehat{M}}_{2}$:
$\max_{x,y}$ $x-y$
subject to $x+y\leq 20,$
$x-y\geq 0,$
$0\leq x\leq 10,$
$0\leq y\leq 5,$
$x,y\in\mathbb{Z}.$
the optimal value of ${\widehat{M}}_{1}$ is ${\widehat{A}}_{1}=10$, while the optimal value of ${\widehat{M}}_{2}$ is ${\widehat{A}}_{2}=0$.
The minor difference in the mathematical models will result in different final values.
Appendix H Examples of Syntax Errors
Some outputs of LLM contain the correct model but not follow the syntax of solvers. For examples, the standard .lp file for
$\max_{x,y}$ $x+y$
subject to $x+y\leq 20,$
$x-y\geq 0,$
$0\leq x\leq 10,$
$0\leq y\leq 5,$
$x,y\in\mathbb{Z}.$
is
Minimize
obj: x + y
Subject To
    c1: x + y <= 20
    c2: x - y >= 0
Bounds
    0 <= x <= 10
    0 <= y <= 5
Generals
    x
    y
End
However, if we replace the constraint $c1$ with x <= 20 - y, the code will be reported as an error, even if the two constraints are equivalent. To specifically test the modelling ability of the LLMs, we apply code modifier to fix the syntax errors.
Appendix I Few-shot Experiment
To test the sensitivity of different models to the number of prompts, we conducted a sensitivity analysis using the scaling law to examine the effect of the number of shots on model performance. In our analysis, we tested the performances of 11 models on ODE problems using prompts with 0 shot, 1 shot, 3 shots, 5 shots, and 10 shots. The results, shown in Figure 6, illustrate the ODE problem-solving accuracy of different models under varying numbers of shots. Similarly, we performed sensitivity tests on optimization problems, assessing the accuracy of 10 models with prompts of 0 shot, 1 shot, 3 shots, and 5 shots. The results are depicted in Figure 7.